OpenStudy (anonymous):

What is 2 and1/4 x 3 and 1/2?

OpenStudy (anonymous):

i think it is 63/8

OpenStudy (anonymous):

2 and 1/4 can be written as 2*(4/4) + 1/4 = 8/4 +1/4 = 9/4 3 and 1/2 can be written as 3*(2/2) + 1/2 = 6/2+1/2 = 7/2 7/2 x 9/4 = 7*9/2*4 = 63/8